OpenStudy (anonymous):

If f(x) = x + 2, find the value of x if f(x) = 12.

OpenStudy (lgbasallote):

12 = x+2 x = 12-2 x = 10

OpenStudy (anonymous):

10

OpenStudy (unklerhaukus):

\[f(x)=x+2\]\[f(x) = 12\] \[x+2=12\]\[x=10\]